Marc Horowitz

24.151

February 8, 1992

(8) ‘English’ is English but ‘French’ is not French.

‘English is English but French’ is not French. ‘English is English but French’ is not ‘French’.

(9) ‘Boring’ is boring and ‘short’ is short but ‘long’ is not ‘short’.

‘Boring’ is ‘boring’ and ‘short’ is short but ‘long’ is not ‘short’. ‘Boring’ is boring and ‘short’ is ‘short’ but ‘long’ is not ‘short’. ‘Boring is boring and short is short but long’ is not short. ‘Boring is boring and short is short but long’ is not ‘short’.

(11) Santa Claus was never born and will never die.

‘Santa Claus’ was never ‘born and will never die’.

Five singular terms:

1. Marc Horowitz

2. The Massachusetts Institute of Technology

3. the CD to which I am listening

4. beeblebrox.mit.edu

5. the set of students registered for 24.151 in Spring, 1992